OpenStudy (anonymous):

how do you do this problem?(7.98x10witha squareroot of -2) x (5.0x10 with a square root -5)

OpenStudy (anonymous):

ok i feel stupid...the -2 and -5 are exponets sorry

OpenStudy (anonymous):

\[7.9x10^{\sqrt{-2}} * 5x10^{\sqrt{-5}}\] \[=39.5*10^{\sqrt{-2}}*10^{\sqrt{-5}}=?\]

OpenStudy (anonymous):

what is x^a *x^b =___?

OpenStudy (anonymous):

x^a *x^b =x^(a+b) can you try 39.5 *10^(sqrt -2)*10^(sqrt-5) =__?